OpenStudy (anonymous):

Using complete sentences, explain how you would use the quadratic formula to solve x2 + 7x = –3. Why is the quadratic formula the best method to use?

myininaya (myininaya):

well first to use the quadratic formula we need ax^2+bx+c=0. so we need to add 3 on both sides such that we have x^2+7x+3=0. now using the quadratic equation, we know a=1, b=7, and c=3. now we just plug those numbers into the quadratic formula which is: \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]. Wherever you see an a, you will put 1. Wherever you see a b, you will put 7. Wherever you see a c, you will put 3. So we have \[x=\frac{-7-\sqrt{7^2-4(1)(3)}}{2(1)}\]. Then I would simply using order of operations. First thing to do evaluate whats in the square root( this is our grouping symbol). The operations tells us to to evaluate the exponent, 7^2=49. Then perform the multiplication as it occurs left from right inside. 4(1)(3)=4(3)=12. Now we do the subtraction inside, we have 49-12=37. so this is what we have now: \[x=\frac{-7 \pm \sqrt{37}}{2(1)}\] so now we can perform the operation on bottom, 2(1)=2. So we have \[x=\frac{-7 \pm \sqrt{37}}{2}\]. This saids we have solutions and they look like \[x=\frac{-7+\sqrt{37}}{2},\frac{7+\sqrt{37}}{2}.\]

myininaya (myininaya):

this saids we have two solutions*